### v1

@@ 6.3 Integrating Syntactic Feedback into Enumerative Search
12:         for $nt\in(prog\times nt^{\prime})$ donormal-$\triangleright$ For all applicable rules
13:              $prog\leftarrow prog.\{nt\rightarrow nt^{\prime}\}$triangleright apply rule to $prog$
14:              $Q\leftarrow Q\cup(c(prog^{\prime})+g(prog^{\prime}),prog^{\prime})$
15:         $i\leftarrow i+1$
@@ 7 Evaluation
We evaluate our approaches on benchmarks taken from the SyGuS competition [3], each with a grammar that corresponds to the full language of their respective theories. We evaluate across three SyGuS categories: Bit-Vector (BV), Linear Integer Arithmetic (LIA), and Invariants (INV). We evaluate both the LLM as a stand-alone synthesizer, the probabilistic enumerator and $A^{*}$ implementations with a pre-trained pCFG and the enumerator with a pre-trained syntactic oracle. We utilize OpenAI’s GPT-3.5-turbo-16k model to generate the prompts used for the pre-trained pCFG and the standalone LLM evaluation, because this model supports longer prompts. We configure this with a temperature of 1.0, conversation-style messaging. We use GPT-3.5-turbo for iLLM-synth, which has shorter prompts. We use the 4.8.12 64-bit version of Z3 for verification and cvc5 version 1.1.0 as a baseline.
@@
We prompt the LLM until it produces up to 6 complete synthesis attempts per benchmark, with the results reported in line 1 of Table 1. Any incomplete solutions are discarded (i.e., functions without a function body), although these are relatively rare, and we discard only $0.85\%$ of programs we generate. In total, the LLM solves $49\%$ of benchmarks, performing better in the invariant and LIA categories than the bitvector category. On average, for the benchmarks it can solve, it takes $4$ attempts to produce a correct solution. The average time for the LLM to generate a program is approximately $5s$ using the OpenAI Python API. However, this is dependent on OpenAI, and we report these times only as estimates in Table 1.
@@
We also report the results obtained by the union of the LLM alone and pCFG-synth, i.e., if the LLM solves the benchmark, we do not deploy the enumerator. This is a more realistic representation of how such a technique would be used and demonstrates that the enumerator can overcome shortcomings of the LLM and vice-versa. The union of the LLM and $A\operatorname*{\!{}^{\ast}}$-pCFG-synth substantially outperforms cvc5, solving $73$ more benchmarks.
@@
Failure Modes
We manually examine a sample of the stand-alone LLM errors and give examples of such errors in Appendix 0.B.1. Broadly, we identify the following common failures: Misunderstandings due to complex constraints (the LLM suggests solutions that are not syntactically close to the correct solution); simple syntactic errors, e.g., applying non-commutative operators to operands in the wrong order, concatenating bitvectors in the wrong order or hallucinating operations; simple semantic errors, e.g., operators in the wrong order. Errors in the first category are not helpful to our guided enumerators, but the remaining categories of error still allow us to generate a wCFG that is likely to indicate the area of the solution. The benchmarks that cvc5 can solve but our enumerative techniques cannot, tend to have complex constraints and relatively short solutions that use less common operators (e.g., bitwise operators). We hypothesize that the LLM guidance becomes an impediment to the enumerator in these scenarios.
@@
$e$-iLLM-synth ${}^{\diamond}$ 241.0 88.2 63.4 9.3 65.3 25.4 370.0 60.8%
$A\operatorname*{\!{}^{\ast}}$-iLLM-synth ${}^{\diamond}$ 272.3 24.6 68.3 20.8 67.3 43.6 408.0 67.0%
${}^{\diamond}$ 142.7 7.2 25.0 1.53 21.0 3.2 188.7 31.0%
@@
Table 1: Summary of results. We run nondeterministic results, marked ${}^{\diamond}$, $3$ times and report the average (standard deviation is less than $1\%$ for all methods except the baseline enumerator for number of benchmarks solved). We highlight the best result in terms of number of benchmarks solved in each category. The timeout is $600$s. Times in italic indicate results that may vary depending on load on the OpenAI servers. The times for pCFG-synth do not include the time to call the standalone LLM and generate the wCFGs, but these are included in the times for LLM $\cup$ pCFG-synth.
@@ 8 Threats to Validity
These experiments use GPT-3.5, an LLM available via API from OpenAI. We have recorded the responses and parameters generated by the LLM in all experiments, but these may not be reproducible [32] since GPT-3.5 behaves non-deterministically in a way that cannot be seeded. However, we observe very small variations in the number of benchmarks solved in our experiments (although greater variation in the average solving time). It is also possible that OpenAI deprecates this LLM and its associated API or updates it and changes its behaviour in the future.
@@ 9 Related Work
Almost all synthesis algorithms use oracles to give feedback to the synthesis process [24, 23]. The majority of these use semantic oracles, which give feedback on the meaning of the program, for example, counterexamples [2]. The LLM in iLLM-synth can be considered a syntactic oracle, giving feedback on the syntax of the program only. Two approaches [16, 1] can be thought of as using syntactic oracles, which evaluate partial programs (or sentential forms) and tell the synthesizer whether a solution can be derived from the sentential form.
@@ 10 Conclusions
We have presented a novel integration of LLMs into two enumerative synthesis algorithms, evaluated on benchmarks from the Syntax-Guided Synthesis competition. We found that LLMs and enumerative solvers have distinct strengths and weaknesses when deployed alone. We have demonstrated that, by allowing the enumerative synthesizer to prompt the LLM with information obtained during the enumeration and allowing the LLM to provide syntactic feedback to the enumeration, we can achieve performance that equals and exceeds the state-of-the-art solvers, even with relatively simple enumerative algorithms. We argue that our results show that LLMs have the potential to make significant contributions in the domain of formal program synthesis but the way to achieve this is by combining these techniques with existing algorithms in the literature. Enumerative synthesis is not dead yet!
Appendix 0.A Initial Ineffective Prompts
In this section, we outline some prompting techniques we tried that did not work.
0.A.0.1 Completing verification conditions:
In the initial phase of our research, we constructed prompts based around asking the LLM to complete an SMT-LIB file that encodes the verification conditions for the synthesis problem, so that it would be unsatisfiable. Consider the following example:
*lollop
(set-logic LIA)
(synth-fun fn0 ((x Int) (y Int)) Int)
(declare-var x Int)
(declare-var y Int)
(constraint (>= (fn0 x y) x))
(constraint (>= (fn0 x y) y))
(constraint (or (= x (fn0 x y)) (= y (fn0 x y))))
(check-synth)
[…]
Example 0: A SyGuS specification that asks for a program that returns $x$ if $x-y\geq 0$ otherwise returns $y$.
The corresponding prompt structure is as follows:
*lollor
; Complete the following SMT file so that it is unsatisfiable
(set-logic LIA)
(declare-fun x () Int)
(declare-fun y () Int)
(assert (or (not (>= (fn0 x y) x))
            (not (>= (fn0 x y) y))
            (not (or (= x (fn0 x y)) (= y (fn0 x y))))))
(check-sat)
(define-fun fn0 ((x Int) (y Int)) Int
[…]
Prompt 0: LLM direct fill-in task.
Here, the LLM’s task was to fill in missing information. This initial approach did not yield satisfactory results.
0.A.0.2 Fine-tuning
[…]
*lollop
[…]
Example 0: Structure for fine-tuning.
Post-adjustment, the Curie model was evaluated using a smaller test set, excluding invariant synthesis, and achieved a 25% success rate in benchmarks. Further refinements to the training set, such as expanding let expressions, did not improve results.
0.A.0.3 Synthesis via Python Code
We experimented with asking the LLM to write solutions in Python and then translate the given solution into an SMT-LIB function.
*lollor
(set-logic LIA)
(synth-fun fn0 ((x Int) (y Int)) Int))
(declare-var x Int)
(declare-var y Int)
(constraint (>= (fn0 x y) x))
(constraint (>= (fn0 x y) y))
(constraint (or (= x (fn0 x y)) (= y (fn0 x y))))
(check-synth)
Write a Python method fn0. Requirements: 1. No built-in Python functions. 2. Adhere to the above SMT-LIB constraints.
[…]
Prompt 0: Prompt for fn0 shown in Example 0.A.0.1.
Upon the LLM generating a possible solution, we employed an additional prompt to convert the Python method into the SMT-LIB function format:
*lollor
Translate the Python method into SMT-LIB code fn0, focusing solely on the SMT-LIB format.
[…]
Prompt 0: Request for converting Python to SMT-LIB code.
This did not work well, as the LLM frequently failed to translate Python into SMT-LIB accurately.
Table 2 presents a comparative study of various prompting methods in solving benchmarks, highlighting the effectiveness of emotional, Lisp, and Python-based prompts. The results indicate a notable enhancement in problem-solving efficiency when using Lisp prompts, with a further improvement observed when emotional stimuli are integrated into the prompts.
Prompt Method Solved %
Lisp 22/50 44%
Emotional 18/50 36%
Python 19/50 38%
Table 2: Comparative analysis of prompt methods: efficacy in solving benchmarks with emotional, Lisp, and Python prompts.
Appendix 0.B Analysis of LLM failures
0.B.1 Unraveling LLM’s Shortcomings
LLM often encounters difficulties with benchmarks characterized by multiple constraints involving longer derivations. These predefined functions contribute to the complexity and length of each constraint, posing a significant challenge for the LLM. For example:
*lollop
(constraint (fn0 (and (= vr0 0) (= 0 vr3)) (= 0 (fn8 vr0 vr3))))
(constraint (fn0 (fn3 (= vr6 (fn8 vr0 vr3)) (>= vr0 vr3) (= vr1 (+ vr0 1)) (= vr7 (+ vr6 1)) (= vr8 (+ vr7 1))) (= vr8 (fn8 vr1 vr3))))
(constraint (fn0 (fn3 (= vr6 (fn8 vr0 vr3)) (>= vr0 vr3) (= vr4 (+ vr3 1)) (= vr7 (+ vr6 1)) (or (and (< vr0 vr4) (= vr9 (+ vr7 1))) (and (>= vr0 vr4) (= vr9 vr7)))) (= vr9 (fn8 vr0 vr4))))
(constraint (fn0 (fn2 (= vr6 (fn8 vr0 vr3)) (< vr0 vr3) (= vr1 (+ vr0 1)) (= vr7 (+ vr6 1))) (= vr7 (fn8 vr1 vr3))))
(constraint (fn0 (fn3 (= vr6 (fn8 vr0 vr3)) (< vr0 vr3) (= vr4 (+ vr3 1)) (= vr7 (+ vr6 1)) (= vr8 (+ vr7 1))) (= vr8 (fn8 vr0 vr4))))
[…]
Example 0: Multiple long constraints for fn8.
If we expand the function applications, we get Example 0.B.1, which still contains long derivations.
*lollop
(constraint (or (not (and (= vr0 0) (= 0 vr3))) (= 0 (fn8 vr0 vr3))))
(constraint (or (not (and (and (and (= vr6 (fn8 vr0 vr3)) (>= vr0 vr3)) (= vr1 (+ vr0 1))) (= vr7 (+ vr6 1))) (= vr8 (+ vr7 1))) (= vr8 (fn8 vr1 vr3))))
(constraint (or (not (and (and (and (= vr6 (fn8 vr0 vr3)) (>= vr0 vr3)) (= vr4 (+ vr3 1))) (= vr7 (+ vr6 1))) (or (and (< vr0 vr4) (= vr9 (+ vr7 1))) (and (>= vr0 vr4) (= vr9 vr7)))) (= vr9 (fn8 vr0 vr4))))
(constraint (or (not (and (and (= vr6 (fn8 vr0 vr3)) (< vr0 vr3)) (= vr1 (+ vr0 1))) (= vr7 (+ vr6 1))) (= vr7 (fn8 vr1 vr3))))
(constraint (or (not (and (and (and (= vr6 (fn8 vr0 vr3)) (< vr0 vr3)) (= vr4 (+ vr3 1))) (= vr7 (+ vr6 1))) (= vr8 (+ vr7 1))) (= vr8 (fn8 vr0 vr4))))
[…]
Example 0: Expanded constraints for fn8.
Such constraints are intricate and challenging for the LLM to understand accurately, making it difficult to produce a response satisfying the specification. The LLM produced a program that contains fn8 itself. However, the pCFG-synth and iLLM-synth can solve the problem with the distribution of rules, whereas the state-of-the-art solver cvc5 cannot.
In contrast, the LLM exhibits a better understanding of benchmarks with shorter derivation for each constraint, such as:
*lollop
(constraint (fn0 (and (< vr1 100) (< vr3 100)) (fn8 vr1 vr2 vr3)))
(constraint (fn0 (fn3 (fn8 vr1 vr2 vr3) (< vr1 100) (< 100 vr3) (fn0 vr4 (and (= vr5 (+ vr1 1)) (= vr6 vr3))) (fn0 (not vr4) (and (= vr5 (- vr1 1)) (= vr6 (- vr3 1))))) (fn8 vr5 vr2 vr6)))
(constraint (fn0 (and (fn8 vr1 vr2 vr3) (not (and (< vr1 100) (< 100 vr3)))) (or (>= vr1 100) (<= vr3 100))))
[…]
Example 0: Shorter derivations in constraints for fn8.
These constraints are less complex, making them more accessible for the LLM to process.
Another challenge arises with benchmarks that necessitate very short solutions. For instance, a benchmark requiring a solution like:
*lollop
(define-fun fn8 ((vr0 Int)) Bool true)
[…]
Program 0: A concise program unattainable by LLM.
In these cases, the LLM tends to propose lengthier solutions, failing to align with the benchmark’s requirement for conciseness. This pattern is evident across all $6$ attempts made by the LLM.
0.B.2 Shuffling Operator or Variable
Below is an example benchmark for finding f:
*lollop
(set-logic BV)
(define-fun hd06 ((x (_ BitVec 32))) (_ BitVec 32)
    (bvor x (bvadd x #x00000001)))
(declare-var x (_ BitVec 32))
(constraint (= (hd06 x) (f x)))
(check-synth)
[…]
Example 0: A SyGuS specification asking for a Bit-Vector program.
LLM failed to provide this exact solution but contributed significantly by suggesting a similar program with the correct operators, albeit in an incorrect order. The program suggested by the LLM is:
*lollop
(define-fun f ((x (_ BitVec 32))) (_ BitVec 32) (bvor x (bvadd x (concat #b1 (_ bv0 31))))).
[…]
Program 0: Correct operators with incorrect order.
In this program, the LLM correctly identifies the use of a bvor and bvadd but misplaces the components of the bit-vector representing the number $1$. The correct solution requires swapping the positions of #b1 and (_ bv0 31) in the concatenation.
This case demonstrates that while the LLM may not always generate a completely accurate solution, it can provide valuable insights or components of a solution. The enumerator, guided by the distribution of rules inferred from the LLM-generated programs, can utilize these insights to reach a correct solution.
0.B.3 Mistaken Conditional Statements
In situations where an individual constructs an if-then-else statement but makes an error in defining the correct condition, we can identify this issue through a specific example of finding fn0:
*lollop
(constraint (>= (fn0 vr0 vr1 vr2 vr3 vr4 vr5 vr6) vr0))
(constraint (>= (fn0 vr0 vr1 vr2 vr3 vr4 vr5 vr6) vr1))
(constraint (>= (fn0 vr0 vr1 vr2 vr3 vr4 vr5 vr6) vr2))
(constraint (>= (fn0 vr0 vr1 vr2 vr3 vr4 vr5 vr6) vr3))
(constraint (>= (fn0 vr0 vr1 vr2 vr3 vr4 vr5 vr6) vr4))
(constraint (>= (fn0 vr0 vr1 vr2 vr3 vr4 vr5 vr6) vr5))
(constraint (>= (fn0 vr0 vr1 vr2 vr3 vr4 vr5 vr6) vr6))
(constraint (or (= vr0 (fn0 vr0 vr1 vr2 vr3 vr4 vr5 vr6)) (or (= vr1 (fn0 vr0 vr1 vr2 vr3 vr4 vr5 vr6)) (or (= vr2 (fn0 vr0 vr1 vr2 vr3 vr4 vr5 vr6)) (or (= vr3 (fn0 vr0 vr1 vr2 vr3 vr4 vr5 vr6)) (or (= vr4 (fn0 vr0 vr1 vr2 vr3 vr4 vr5 vr6)) (or (= vr5 (fn0 vr0 vr1 vr2 vr3 vr4 vr5 vr6)) (= vr6 (fn0 vr0 vr1 vr2 vr3 vr4 vr5 vr6)))))))))
[…]
Example 0: SyGuS specification for constructing a program contains several $if\text{-}then\text{-}else$ Operators.
In this scenario, the accurate and appropriate solution should be
*lollop
(define-fun fn0 ((vr0 Int) (vr1 Int) (vr2 Int) (vr3 Int) (vr4 Int) (vr5 Int) (vr6 Int)) Int
    (ite (and (>= vr0 vr1) (>= vr0 vr2) (>= vr0 vr3) (>= vr0 vr4) (>= vr0 vr5) (>= vr0 vr6))
[…]
Program 0: Solution program for Example 0.B.3.
However, the LLM generates:
*lollop
(define-fun fn0 ((vr0 Int) (vr1 Int) (vr2 Int) (vr3 Int) (vr4 Int) (vr5 Int) (vr6 Int)) Int
   (ite (and (<= vr0 vr1) (<= vr1 vr2) (<= vr2 vr3) (<= vr3 vr4) (<= vr4 vr5) (<= vr5 vr6))
[…]
Program 0: Program generated by LLM for Example 0.B.3.
, which, although sharing a similar structure, inaccurately defines the conditions.
0.B.4 Unorthodox Nesting and Syntactically Incorrect
A common problem occurs when LLM endeavors to construct a nested if-then-else statement. However, their nesting level is insufficient, and they also introduce operators incompatible with SyGuS.
For instance, while the correct solution should involve a properly nested if-then-else structure
*lollop
(define-fun fn19 ((vr0 Int) (vr1 Int)) Int
(ite (and (>= (+ vr0 vr1) 2)
    (ite (>= (+ vr0 vr1) 3)
        (ite (>= (+ vr0 vr1) 4)
            (ite (>= (+ vr0 vr1) 5)
                (or (>= (+ vr0 vr1) 6)
                    (= vr0 (+ 11 (* (- 1) vr1))))
                (= vr0 (+ 1 (* (- 1) vr1))))
            (= vr0 (+ 3 (* (- 1) vr1))))
        (= vr0 (+ 1 (* (- 1) vr1)))))
    (+ (- 60) (* 60 vr0) (* 60 vr1))
    (ite (ite (>= (+ vr0 vr1) 2)
        (and (>= (+ vr0 vr1) 3)
            (ite (>= (+ vr0 vr1) 4)
                (ite (>= (+ vr0 vr1) 5)
                    (or (= vr0 (+ 11 (* (- 1) vr1)))
                        (not (>= (+ vr0 vr1) 6)))
                    (= vr0 (+ (- 9) (* (- 1) vr1))))
                (= vr0 (+ (- 1) (* (- 1) vr1)))))
        (= vr0 (+ (- 1) (* (- 1) vr1))))
        (+ 50 (* 50 vr0) (* 50 vr1))
        (ite (and (>= (+ vr0 vr1) 2)
            (ite (>= (+ vr0 vr1) 3)
                (ite (>= (+ vr0 vr1) 4)
                    (or (not (>= (+ vr0 vr1) 5))
                        (ite (>= (+ vr0 vr1) 6)
                            (= vr0 (+ 1 (* (- 1) vr1)))
                            (= vr0 (+ (- 9) (* (- 1) vr1)))))
                    (= vr0 (+ 7 (* (- 1) vr1))))
                (= vr0 (+ 1 (* (- 1) vr1)))))
            (+ (- 40) (* 40 vr0) (* 40 vr1))
            (ite (and (>= (+ vr0 vr1) 2)
                (ite (>= (+ vr0 vr1) 3)
                    (ite (>= (+ vr0 vr1) 4)
                        (or (>= (+ vr0 vr1) 6)
                            (not (>= (+ vr0 vr1) 5))
                            (not (= vr0 (+ (- 1) (* (- 1) vr1)))))
                        (not (= vr0 (+ (- 1) (* (- 1) vr1)))))
                    (not (= vr0 (+ 3 (* (- 1) vr1))))))
                (+ (- 20) (* 20 vr0) (* 20 vr1))
                (+ 10 (* 10 vr0) (* 10 vr1))))))
[…]
Program 0: Solution program with nested $if\text{-}then\text{-}else$ operators.
The LLM generates the program below Program 0.B.4, which does not achieve the required nesting level and erroneously incorporates operators. For example, if is syntactically incorrect in the SyGuS framework.
*lollop
(define-fun fn19 ((vr0 Int) (vr1 Int)) Int
    (ite (<= (fn1 vr0 vr1) 1)
        (if (= (fn1 vr0 vr1) 1)
            (fn17 (fn2 vr0 vr1 1))
            0)
        (ite (<= (fn1 vr0 vr1) 2)
            (if (= (fn1 vr0 vr1) 2)
                (fn17 (fn10 (fn2 vr0 vr1 (- 1))))
                0)
            (ite (<= (fn1 vr0 vr1) 3)
                (if (= (fn1 vr0 vr1) 3)
                    (fn17 (fn11 (fn2 vr0 vr1 1)))
                    0)
                (ite (<= (fn1 vr0 vr1) 4)
                    (if (= (fn1 vr0 vr1) 4)
                        (fn17 (fn12 (fn2 vr0 vr1 (- 1))))
                        0)
                    (ite (<= (fn1 vr0 vr1) 5)
                        (if (= (fn1 vr0 vr1) 5)
                            (fn17 (fn13 (fn2 vr0 vr1 1)))
                            0)
                        (fn17 (fn14 (fn2 vr0 vr1 (- 1))))))))))
[…]
Program 0: Program with insufficient nesting and syntactic incompatibilities.

### v2

@@ 6.3 Integrating Syntactic Feedback into Enumerative Search
12:         for $(nt\in prog)\times nt^{\prime}$ do
13:              if $(nt\times nt^{\prime})\in P_{G}.R$ then $\triangleright$ For all applicable rules
14:                  $prog\leftarrow prog.\{nt\rightarrow nt^{\prime}\}$triangleright apply rule to $prog$
15:                  $Q\leftarrow Q\cup(c(prog)+g(prog),prog)$
16:         $i\leftarrow i+1$
@@ 7 Evaluation
We evaluate our approaches on benchmarks taken from the SyGuS competition [3], each with a grammar that corresponds to the full language of their respective theories. We evaluate across three SyGuS categories: Bit-Vector (BV), Linear Integer Arithmetic (LIA), and Invariants (INV). We evaluate both the LLM as a stand-alone synthesizer, the probabilistic enumerator and $A^{*}$ implementations with a pre-trained pCFG and the enumerator with a pre-trained syntactic oracle. We utilize OpenAI’s GPT-3.5-turbo-16k model to generate the prompts used for the pre-trained pCFG and the standalone LLM evaluation because this model supports longer prompts. We configure this with a temperature of 1.0, conversation-style messaging. We use GPT-3.5-turbo for iLLM-synth, which has shorter prompts. We use the 4.8.12 64-bit version of Z3 for verification and cvc5 version 1.1.0 as a baseline.
@@
We prompt the LLM until it produces up to 6 complete synthesis attempts per benchmark, with the results reported in line 1 of Table 1. Any incomplete solutions are discarded (i.e., functions without a function body), although these are relatively rare, and we discard only $0.85\%$ of programs we generate. In total, the LLM solves $49\%$ of benchmarks, performing better in the invariant and LIA categories than the bit-vector category. On average, for the benchmarks it can solve, it takes $4$ attempts to produce a correct solution. The average time for the LLM to generate a program is approximately $5s$ using the OpenAI Python API. However, this is dependent on OpenAI, and we report these times only as estimates in Table 1. We allow the LLM only $6$ attempts to solve the problem since, by the $6^{th}$ iteration, the number of new solutions the LLM finds has dropped to $<2\%$ (and it finds 00 new solutions for LIA).
@@
We also report the results obtained by the union of the LLM alone and pCFG-synth, i.e., if the LLM solves the benchmark, we do not deploy the enumerator. This is a more realistic representation of how such a technique would be used and demonstrates that the enumerator can overcome shortcomings of the LLM and vice versa. The union of the LLM and $A\operatorname*{\!{}^{\ast}}$-pCFG-synth substantially outperforms cvc5, solving $73$ more benchmarks.
@@
Failure Modes:
We manually examine a sample of the stand-alone LLM errors and give examples of such errors in the extended version of this paper  [^2] https://arxiv.org/html/2403.03997. Broadly, we identify the following common failures: Misunderstandings due to complex constraints (the LLM suggests solutions that are not syntactically close to the correct solution); simple syntactic errors, e.g., applying non-commutative operators to operands in the wrong order, concatenating bit-vectors in the wrong order or hallucinating operations; simple semantic errors, e.g., operators in the wrong order. Errors in the first category are not helpful to our guided enumerators, but the remaining categories of error still allow us to generate a wCFG that is likely to indicate the area of the solution. The benchmarks that cvc5 can solve and our enumerative techniques cannot, tend to have complex constraints and relatively short solutions that use less common operators (e.g., bitwise operators). We hypothesize that the LLM guidance becomes an impediment to the enumerator in these scenarios. In contrast, the average length (in characters) of a solution for benchmarks uniquely solved by the LLM is 4.7x the length of a solution for benchmarks uniquely solved by cvc5. Using the LLM to guide the enumerators increases the length of solutions that the enumerators can find, for instance all solutions found by $A^{*}$ contain fewer than $3$ operators, but $A^{*}$-iLLM-synth finds solutions with greater than $20$ operators.
Programming-by-example:
We omit benchmarks from the syntax-guided synthesis competition tracks that solely focus on programming-by-example (PBE)(i.e., specifying a program only using input-output examples and a grammar). We omit these benchmarks for two reasons: first, since training data is trivial to generate for PBE, unlike general logical specifications [34], there are many other successful machine-learning driven synthesis techniques that can be trained for PBE techniques[6]. Second; our approaches are effective when the LLM can provide guidance to the enumerator, which comes from prompting the LLM with the logical constraints that form the specification. If we prompt the LLM using the prompting techniques outlined in Section 4.1 with a PBE specification, it tends to provide a solution in the form of a large case split over the input examples, which returns specific outputs for each input. This is not useful for guiding the enumerator because the LLM overfits to the examples in the specification and fails to provide any bias towards operators other than “if-then-else”. To extend our approach to PBE, we would need to use a prompting approach tailored to input-output examples.
@@
$e$-iLLM-synth 241.0 88.2 63.4 9.3 65.3 25.4 370.0 60.8%
$A\operatorname*{\!{}^{\ast}}$-iLLM-synth 272.3 24.6 68.3 20.8 67.3 43.6 408.0 67.0%
enumerator 142.7 7.2 25.0 1.53 21.0 3.2 188.7 31.0%
@@
Table 1: Summary of results. We run nondeterministic results, marked , $3$ times and report the average (standard deviation is less than $1\%$ for all methods except the baseline enumerator for number of benchmarks solved). We highlight the best result in terms of number of benchmarks solved in each category. The timeout is $600$s. Times in italic indicate results that may vary depending on load on the OpenAI servers. The times for pCFG-synth do not include the time to call the standalone LLM and generate the wCFGs, but these are included in the times for LLM $\cup$ pCFG-synth.
@@ 8 Threats to Validity
These experiments use GPT-3.5, an LLM available via API from OpenAI. We have recorded the responses and parameters generated by the LLM in all experiments, but these may not be reproducible [33] since GPT-3.5 behaves non-deterministically in a way that cannot be seeded. However, we observe very small variations in the number of benchmarks solved in our experiments (although greater variation in the average solving time). It is also possible that OpenAI deprecates this LLM and its associated API or updates it and changes its behavior in the future.
@@ 9 Related Work
Almost all synthesis algorithms use oracles to give feedback to the synthesis process [25, 24]. The majority of these use semantic oracles, which give feedback on the meaning of the program, for example, counterexamples [2]. The LLM in iLLM-synth can be considered a syntactic oracle as it only gives feedback on the syntax of the program. Two approaches [17, 1] can be thought of as using syntactic oracles, which evaluate partial programs (or sentential forms) and tell the synthesizer whether a solution can be derived from the sentential form.
@@ 10 Conclusions
We have presented a novel integration of LLMs into two enumerative synthesis algorithms, evaluated on benchmarks from the Syntax-Guided Synthesis competition. We found that LLMs and enumerative solvers have distinct strengths and weaknesses when deployed alone. We have demonstrated that, by allowing the enumerative synthesizer to prompt the LLM with information obtained during the enumeration and allowing the LLM to provide syntactic feedback to the enumeration, we can achieve performance that equals and exceeds the state-of-the-art solvers, even with relatively simple enumerative algorithms. We argue that our results show that LLMs have the potential to make significant contributions in the domain of formal program synthesis, but the way to achieve this is by combining these techniques with existing algorithms in the literature. Enumerative synthesis is not dead yet!
Acknowledgements:
This work was in part supported by an Amazon Research Award, a Royal Academy of Engineering research fellowship, and the European Research Council (ERC) project FormalWeb3 (Grant ID 101156734).